A number consists of two digits whose sum is 8. If 18 is added to the number its digit are reversed. Find the number.
Answer by anand429(138)

Let digits be x and y
Number = 10x+y
Reversed Number = 10y+x
As per ques.
x+y = 8 --- (i)
10x+y +18 = 10y+x
=> 9x-9y = -18
=> x-y = -2 ----(ii)
Adding (i) and (ii)
2x =6
x=3
so y =5
so number is 35